OpenStudy (anonymous):

Perform the indicated opperation as a fraction in lowest terms Question: 21/50 - 9/14 ?

jimthompson5910 (jim_thompson5910):

\[\Large \frac{21}{50}-\frac{9}{14 }\] \[\Large \frac{21}{50}\times \frac{7}{7}-\frac{9}{14 }\times \frac{25}{25}\] \[\Large \frac{21 \times7}{350}-\frac{9 \times25}{350}\] \[\Large \frac{147}{350}-\frac{225}{350}\] \[\Large \frac{147-225}{350}\] \[\Large \frac{-78}{350}\] \[\Large -\frac{39}{175}\] So \[\Large \frac{21}{50}-\frac{9}{14 } = -\frac{39}{175}\]

jimthompson5910 (jim_thompson5910):

The basic idea is to get a common denominator. Once you get the LCD, you can combine the fractions.